OpenStudy (anonymous):

Rohan went to the market to buy 10 kg of each of oranges, mangoes, bananas and grapes. The cost of 5 kg of oranges and 2 kg of mangoes together was rs 310. The cost of 3 kg of mangoes and 3.5 kg of bananas together was rs 230. The cost of 1.5 kg of bananas and 5 kg of grapes together was rs 160. Find the amount spent by Rohan.

OpenStudy (anonymous):

3.5kg banana=230 and 1.5kg banana=160 so 5kg banana =230+160=390 3kg mangoes=230,2kg mangoes=310 so 5kg mangoes=230+310=540 5 kg grapes=160 5kg pranges=310

OpenStudy (anonymous):

now add up all these 390+540+160+310=1400

OpenStudy (anonymous):

the answer is correct. but i see that 390+540+160+310= 1400(which is 40kgs in total.) but i don't understand the reasoning behind 3kg of mangoes + 3.5kg of bananas = 230. when you have taken simply 3.5kg of banana as 230 and 3kg of mango as 230. can you explain the approach to this? what am i missing?

OpenStudy (anonymous):

3 kg mangoes and 3.5 kg bananas collectively make 230. how are you isolating them to determine the price?